A magazine reported the results of a random telephone poll commissioned by a television network. Of the1209 men who​ responded, only 20 said that their most important measure of success was their work. Complete parts a through c. a right parenthesis a) Estimate the percentage of all males who measure success primarily from their work. Use a 99​% confidence interval.​ Don't forget to check the conditions first.
Answer #1

Let , X be the number of males who measures success primarily from their work.

Given : n=1209 , x=20

a) The estimate of the percentage of all males who measures success primarily from their work. is ,

p=x/n=20/1209=0.0165

Conditions :

1) The data is simple random sample from the normal distribution.

2) The sample values are independent

3) np=1209*0.0165=20>10 and nq=1209*(1-0.0165)=1209*0.9835=1189.0515>10

b) Therefore , the 99% confidence interval for proportion is ,